[Hallc_running] [New Logentry] RC Update Wed Feb 26

brads at jlab.org brads at jlab.org
Wed Feb 26 17:50:04 EST 2020

Logentry Text:
<p><strong>https://hallcweb.jlab.org/wiki/index.php/RC_Meeting_and_Run_Plan,_Tuesday,_February_26th_2020<br></strong></p><p><strong>Wednesday 26th Feb. 2020. </strong></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><ul><li>Beam back around 7pm (4 hour MCC 'maintenance down' on Tuesday => ~11 hours)<ul><li>Ran production (longitudinal) overnight, a beam a bit trippy</li></ul></li><li>MCC/OPS safety stand down 9--12pm; beam back around 12:45pm, more or less on schedule!<ul><li>JP did Target transverse polarization tests this morning during the Safety stand down. See log entry <a class="external free" href="https://logbooks.jlab.org/entry/3792457" rel="nofollow">https://logbooks.jlab.org/entry/3792457</a></li></ul></li></ul><ul><li>Moller cooldown completed<ul><li>Moller measurement Thursday day</li></ul></li></ul><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li>Injector spot move tentatively pl!
 anned for early next week (Mon/Tues)?<ul><li>2--3 hour down</li></ul></li><li>Photocathode reactivation likely deferred until week of March 16(?)<ul><li>Will depend on measured QE lifetime after spot move</li></ul></li></ul><ul><li>Target coil rotation planned for March 13--20<ul><li>Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program</li><li>Lots of preplanning to be completed (RadCon, Techs, S&A, Target group, etc)</li><li>RW-II training for JP, Arun, Brad (Complete the training documentation and take the RW-II test if you haven't already done so)</li></ul></li></ul><h2><span id="Pending_problems" class="mw-headline">Pending problems</span></h2><ul><li>Ongoing target correction coil readback/software issue (LabView) <a class="external free" href="https://logbooks.jlab.org/entry/3792556" rel="nofollow">https://logbooks.jlab.org/entry/3792556</a></li><li>Persistent BLM & ion chamber trips<ul><li>General activatio!
 n, weird BLM issues (<a class="external free" href="https://lo!
 gbooks.jlab.org/entry/3792326" rel="nofollow">https://logbooks.jlab.org/entry/3792326</a>), Ad-hoc ion chamber threshold adjustments (<a class="external free" href="https://logbooks.jlab.org/entry/3792095" rel="nofollow">https://logbooks.jlab.org/entry/3792095</a>)</li><li><strong>NOTE</strong>: Shift crew should <em>NOT</em> give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)</li></ul></li></ul><h2><span id="Fixed_problems" class="mw-headline">Fixed problems</span></h2><h2><span id="Opportunistic_Access_Jobs" class="mw-headline">Opportunistic Access Jobs</span></h2><ul><li>Target group would like to check NMR connections and hardware settings in the Hall (JP, Junhao)</li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3792679
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200226/b40db9c7/attachment.html>

More information about the Hallc_running mailing list